What is the molality of 14 g of MgSO4 in 250 mL of water at 40...

  1. What is the molality of 14 g of MgSO4 in 250 mL of water at 40 °C?
Answer #1

density of water at 40C = 0.992 g/ml

mass of water = volume * density

= 250 * 0.992 = 248 grams or 0.248 Kg

moles of mgso4 = mass / molar mass

= 14 / 120.37

= 0.11631

molality = moles / mass of solvent (in kg)

= 0.11631 / 0.248

= 0.469

any query please comment

if satisfied please rate it thanks